Calculate the amount of heat (in kJ) required to raise the temperature of a 88.0 g sample of ethanol from 298.0 K to 405.0 K. The specific heat capacity of ethanol is 2.42 J/g⋅∘C.
Answer #1

Given mass of ethanol m = 88 g

Initial temperature of ethanol T1 = 298 K = (298-273)= 25℃

Final temperature of ethanol T2= 405 K = (405-273) = 132℃

heat capacity of ethanol, c = 2.42 J/g ℃

the heat (Q) required is given as:

Q = mcΔT = mc(T2-T1)

Q = 88.0 g × 2.42 J/g℃×(132-25) C = 22786 J

So Heat required = 22.79 kJ

Thank you.